OpenStudy (anonymous):

A circus car with a clown has a total mass of 150 kg. It is coasting at 6.0 m/, when it hits a large spring head on. If it is brought to a stop the time the spring is compressed 2.0 m, what is the force constant of spring?

OpenStudy (anonymous):

kinetic Energy of car will be stored as P.E of Spring \[\frac{ 1 }{ 2 }mv^2 = \frac{ 1 }{ 2 }kx^2\] \[150*6*6 = k*2*2\] \[k=1350N/m\]

OpenStudy (anonymous):

thanks a lot

OpenStudy (anonymous):

Welcome..)